A STUDY OF 2H- AND 2D-BENZOTRIAZOLE IN THEIR LOWEST ELECTRONIC STATES BY UV-LASER DOUBLE RESONANCE SPECTROSCOPY

WOLFGANG ROTH, CHRISTOPH JACOBY, ARNIM WESTPHAL AND \underline{MICHAEL SCHMITT}, Institut für Physikalische Chemie und Elektrochemie I, Heinrich-Heine-Universität Düsseldorf, 40225 Düsseldorf, Germany.

Hole burning spectra of 2H -benzotriazole and 2D- benzotriazole have been recorded in a range of 1250 cm^{−1} above the electronic origin to rule out the possibilty that the other tautomer (1H -benzotriazole) absorbs in this frequency range.^{a} The 1H -conformer could be detected via FTIR in the region of the NH-stretch vibration and distinguished from the 2H -conformer by its different band contour. The relative stabilities of both conformers have been determined from a simulation of the temperature dependent amount of each conformer and compared to the results of ab initio calculations. After establishing the existence of only one absorbing species by spectral hole burning, dispersed fluorescence spectra taken through several low frequency vibronic bands of 2H (2D)-benzotriazole have been recorded. The ground state vibrational frequencies are compared to the results of an ab initio MP2 based normal mode analysis. Assignments of S_{1} vibrations to specific ground state vibrations were established by comparison to calculations. The rotational band contours of the vibronic 2H -benzotriazole transitions have been determined to be pure a- or b-type, while rotationally resolved LIF showed the electronic origin of benzotriazole to be pure b-type.^{b} This can be traced back to the existence of two close lying electronically excited states of 2H -benzotriazole. While the electronic origin, investigated by Berden et al.^{b} can be attributed to an ^{1}B_{2}(^{1}L_{a}) ← ^{1}A_{1} transition, the a- type transitions end up in the ^{1}L_{b} state. The large number of vibronic levels and the possible mixing of states prevents a clear assignment of the observed transitions in the S_{1}- and S_{2}-state.

aW. Roth, Ch. Jacoby, A. Westphal and M. Schmitt accepted for publication in J. Phys. Chem. (1989)

bG. Berden, E. Jalviste and W.L. Meerts Chem. Phys. Lett. 226 (1994) 305-309